OpenStudy (anonymous):

The product of a and b is x , and the sum of a and b is y If x and y are both positive numbers, what is 1/a + 1/b in terms of x and y?

OpenStudy (anonymous):

y/x

OpenStudy (anonymous):

1/a+1/b=(a+b)/ab=y/x

OpenStudy (kira_yamato):

\[y/x\]

OpenStudy (anonymous):

1/a + 1/b = (a+b)/(ab) = y/x.